# Joint Channel, Data and Radar Parameter Estimation for AFDM Systems in Doubly-Dispersive Channels

### v2

@@ I Introduction
The building blocks of communications and radar systems have been traditionally designed separately and assumed to utilize orthogonal resources. Integrated sensing and communications (ISAC)  [^1] Also commonly referred to as joint communication and sensing (JCAS) and joint communications and radar (JCR). [2, 3, 4], a recently-emerged technology seen by many as a key driver of future mobile communication systems, challenges that established concept by considering the simultaneous provision of radar and communications functionalities under the same architecture, having the advantages of more efficient use of resources, lower power consumption, decreased hardware costs and the potential for wide market penetration [5, 6, 7].
Staying true to the integrative aspect of the proposition [^2] We distinguish ISAC from earlier co-existence approaches [21, 22], without integration of these functions., earlier contributions on ISAC focused on schemes employing existing communications waveforms. In particular, many methods found in current literature seek to take advantage of certain features of channel estimation – especially in higher frequency bands where less scattering and higher resolution enable the estimation of individual channel paths – to incorporate radar-like functionalities, such as the estimation of radar parameters from echoes of the transmit signal [8, 9, 10, 11, 12].
Examples of the latter are techniques based on the orthogonal frequency division multiplexing (OFDM) waveform [10, 11] which, however, indicated that the high inter-carrier interference (ICI) can be a bottleneck that reduces robustness to high Doppler shifts. This factor alone, which stems from the fact that the channel diagonals spread into a decaying band [13], severely hinders the use of OFDM in high mobility scenarios [14], as envisioned for 6G systems.
Another potential candidate for the ISAC paradigm is the orthogonal time frequency space (OTFS) waveform, which was originally proposed in [15] as a novel two-dimensional (2D) modulation scheme that embeds information directly on the delay-Doppler (DD) domain. Since the DD domain had a direct parallel to common radar processing framework, OTFS was quite popular at its inception, attracting much work that demonstrated its effectiveness for ISAC [8, 9, 10, 12, 11]. Besides the significantly higher complexity compared to OFDM, which is inherent to its 2D modulation approach, another inherent limitation of OTFS is that it does not achieve the optimal diversity order in doubly-dispersive channels [16].
@@
It has been in fact demonstrated that AFDM has a superior bit error rate (BER) performance compared to OFDM, achieving also BERs comparable to OTFS while requiring lower overhead for channel estimation [19]. In turn, evidence that AFDM can be effectively utilized for ISAC has also been given, for instance in [18] where a simple matched-filter (MF) approach was suggested. Building on the latter, an improved AFDM-base ISAC performance was also achieved in [20], albeit at the expense of a significant increase in computational complexity due to the exhaustive maximum likelihood (ML) grid search required.
@@
One aspect of the ISAC concept that was often overlooked, and which recently has been addressed more frontally, however, is the impact of pilot overhead onto the overall approach. It has been shown in [19], for instance, that a major advantage of AFDM over OFDM and OTFS is the lighter requirement in the number of pilots, which in turn can be taken advantage of by enabling the base station (BS) to perform radar parameter estimation (RPE) without expensive self-interference cancellation (SIC) techniques. The trend to reduce pilot overhead suggests, therefore, the adoption of joint channel and data estimation (JCDE) technology, which has been shown to be very effective in allowing efficient channel estimation with very few pilots [^3] The embedded channel estimation method proposed in [19] requires a single pilot, but surrounded by a good amount of zeros, such that the fraction of the AFDM frame dedicated to payload is affected not just by the sole pilot but also the zero-padding. and under strenuous conditions [23, 24, 25, 26, 27, 1], significantly reduce the overhead of coherent wireless systems. One example of the latter is the Bayesian bilinear inference (BBI) JCDE scheme for OFDM systems proposed [27], which was shown to have the ability of handling channel aging and to entirely remove the need of pilots in favor of adequate coding of an equivalent rate. Another is the work in [1], where a Bayesian parametric bilinear Gaussian belief propagation (PBiGaBP) JCDE scheme for OTFS systems was proposed, which was shown to asymptotically approach the idealized scheme where perfect channel state information (CSI) knowledge is available.
@@
In view of all the above, we propose in this paper a combined and generalized JCDE and ISAC solution for systems known to cope well with doubly-dispersive channels, including OFDM, OTFS and AFDM. The proposed solution comprises two main contributions. The first is a new JCDE, designed with basis on the Bayesian PBiGaBP originally proposed for OTFS waveforms in [1], which is shown to function also for the other modulation schemes under the general doubly dispersive channel model described in [13]. Comparisons of the state-of-the-art (SotA) OTFS and the contributed AFDM methods against a genie-aided linear Gaussian belief propagation (GaBP) benchmark scheme, in which perfect CSI is assumed, demonstrate that the proposed solution under AFDM achieves the same performance of the OTFS SotA, both of which also closely approach the idealistic benchmark, indicating that there is no penalty for the reduced modulation complexity advantage obtained when utilizing AFDM instead of OTFS, and that the proposed scheme is effective in reducing pilot overhead.
The second contribution is a new RPE scheme based on the vectorized probabilistic data association (PDA) approach [24, 28] incorporating a Bernoulli-Gaussian denoiser tuned via expectation maximization (EM) [29] which quickly converges to the appropriate distribution parameters.
@@
Compared to the SotA alternative of [30], which is designed around the sparse Bayesian learning (SBL) algorithm, the contributed method is shown to outperform the latter over any of the relevant waveforms, namely, OFDM, OTFS in [30] and AFDM. The technique also has an inherent and substantial computational cost advantage over ML-based SotA alternatives such as those in [19] and [20].
@@
  • In Section II, we concisely revise the unified OFDM, OTFS and AFDM system model for doubly-dispersive channels from [13].
  • Taking advantage of the latter, we propose in Section III a PBiGaBP-based scheme for JCDE, which applies to all the aforementioned waveforms.
  • We show that even if a single pilot per block is used, the proposed JCDE-AFDM system significantly outperforms both the OFDM and OTFS alternatives [1], as well as the AFDM method in [19].
  • Finally, in Section IV-B a novel PDA-based method for RPE in doubly-dispersive channels is proposed, which in addition to enabling monostatic implementation without requiring full-duplex (FD) technology, is shown to outperform the SotA of [30].
@@ II System Model
We follow [20] and consider an ISAC scenario composed of a single-antenna [^4] The generalization of the proposed methods to the multiple-input multiple-output (MIMO) case requires the formulation and solution of a non-trivial transmit (TX) and receive (RX) beamforming optimization problem, with objective and constraints designed to satisfy both communication and sensing requirements, but which otherwise have no fundamental influence on the JCDE and RPE schemes that will be proposed. That contribution is therefore out of the scope of this article and will be offered in a follow-up work. BS transmitting in downlink and taking advantage of first-order reflected (echo) signals to perform RPE in order to detect the ranges and velocities of objects (targets) in its surroundings, while the user equipment (UE) performs JCDE as illustrated in Figure 1. As depicted in the figure, it is assumed that a total of $P_{\mathrm{B}}$ target-originated echo signals reach the BS, which might be either from UE or other objects in the ambient. In turn, each UE is assumed to receive the signal transmitted by the BS scattered onto $P_{\mathrm{U}}+1$ paths, with $p_{\mathrm{U}}=0$ denoting the line-of-sight (LoS), without loss of generality (wlg), and the remaining $P_{\mathrm{U}}\leq P_{\mathrm{B}}$ paths being non-LoS (NLoS) components of the multi-path channel.
For the sake of simplicity and again wlg, we shall hereafter attain to a single user, since the multi-user (MU) case has no impact onto the JCDE procedure performed independently at each user, and is transparent to the RPE algorithm, as long as the radar receiver has sufficient degrees-of-freedom (DoF) to resolve all $P_{\mathrm{B}}$ echoed signals it receives. Also for simplicity, although we differentiate $P_{\mathrm{B}}$ and $P_{\mathrm{U}}$ in Fig. 1, in the equations that follow, the number of paths/targets will be denoted by $P$, with the distinction between the two cases left to context.
$\mathbf{\Phi}_{p}\triangleq\text{diag}\bigg{(}[\overbrace{e^{-j2\ell_{p}\pi},e^{-j2(\ell_{p}-1)\pi},\dots,e^{-j4\pi},e^{-j2\pi}}^{\ell_{p}\;\text{terms}},\overbrace{1,1,\dots,1,1}^{N-\ell_{p}\;\text{ones}}]\bigg{)}\in\mathbb{C}^{N\times N}.\vspace{-2ex}$ (5)
@@
Figure 1: Illustration of an ISAC system with $P_{\mathrm{B}}$ target echoes received by the BS, and $P_{\mathrm{U}}$ resolvable paths seen at the UE.
II-A Generalized Doubly-Dispersive Channel Model Model
@@
where $t$ and $\tau$ are the continuous time and delay variables, respectively.
Leveraging equation (1), the input-output relationship of the channel in the TD can be obtained via the linear convolution of the transmit signal $s(t)$ and channel impulse response as described in [33, eq. (10.69)], which added with noise yields
$r(t)=s(t)*h(t,\tau)+w(t)=$
@@
Defining $r[n]$ and $s[n]$, with $n\in\{0,\dots,N-1\}$, to be the sampled sequences of $r(t)$ and $s(t)$, respectively, with samples taken at the rate $f_{\mathrm{S}}\triangleq\frac{1}{T_{\mathrm{S}}}$[Hz], the following discrete equivalent of equation (II-A) can be obtained
$r[n]=\sum_{\ell=0}^{\infty}s[n-\ell]\bigg{(}\sum_{p=0}^{P}h_{p}\cdot e^{j2\pi\frac{\nu_{p}}{f_{\mathrm{S}}}n}\cdot\delta\Big{(}\ell-\frac{\tau_{p}}{T_{\mathrm{S}}}\Big{)}\bigg{)}+w[n],$ (3)
where $\ell$ is the discrete delay index, and $T_{\mathrm{S}}$ the delay resolution as well as the sampling interval, such that $\tau=\ell\cdot T_{\mathrm{S}}$.
Finally, defining the normalized digital Doppler shift and the normalized delay of the $p$-th path respectively as $f_{p}\triangleq\frac{N\nu_{p}}{f_{\mathrm{S}}}$ and $\ell_{p}\triangleq\frac{\tau_{p}}{T_{\mathrm{S}}}$, where the delay resolution $T_{\mathrm{S}}$ is assumed to be sufficiently high to ensure that $\ell_{p}-\lfloor\frac{\tau_{p}}{T_{\mathrm{S}}}\rceil\approx 0$, the discrete convolution described in equation (3) can be rewritten in terms of the circular convolution
$\mathbf{r}=\bigg{(}\sum_{p=0}^{P}h_{p}\cdot\mathbf{\Phi}_{p}^{\phi_{\text{CP}}}\cdot\bm{\Omega}^{f_{p}}\cdot\mathbf{\Pi}^{\ell_{p}}\bigg{)}\cdot\mathbf{s}+\mathbf{w}=\mathbf{H}\cdot\mathbf{s}+\mathbf{w},\vspace{-0.5ex}$ (4)
where the inclusion and posterior removal of a cyclic prefix (CP) of length $N_{\text{CP}}$ to the sequence $s[n]$ is already taken into account [13].
In equation (4), $\mathbf{r}\triangleq[r[0],\cdots,r[N-1]]\in\mathbb{C}^{N\times 1}$ and $\mathbf{s}\triangleq[s[0],\cdots,s[N-1]]\in\mathbb{C}^{N\times 1}$ are the transmit and received signal sample vectors; $\mathbf{w}\in\mathbb{C}^{N\times 1}$ denotes AWGN; $\mathbf{H}\in\mathbb{C}^{N\times N}$ is the effective channel matrix; $\mathbf{\Phi}_{p}\in\mathbb{C}^{N\times N}$, described in equation (5), is such that $\mathbf{\Phi}_{p}^{{}_{\text{CP}}}$ captures the result of the inclusion of the CP onto the $p$-th channel path; $\bm{\Omega}\in\mathbb{C}^{N\times N}$, described in equation (6), is a diagonal matrix containing the $N$ complex roots of one and $\mathbf{\Pi}\in\{0,1\}^{N\times N}$ is the forward cyclic shift matrix [^6] To clarify, the matrix $\mathbf{\Pi}$ is such that $\bm{A}\cdot\mathbf{\Pi}^{\ell_{p}}$, with $\ell_{p}\in\mathbb{N}_{0}$, is a cyclic left-shifted version of the matrix $\bm{A}$, such that the first $\ell_{p}$ columns of $\bm{A}$ are moved to the positions of the last $\ell_{p}$ columns. Notice also that $\mathbf{\Pi}^{0}$ is the $N\times N$ identity matrix, such that $\bm{A}\cdot\mathbf{\Pi}^{0}=\bm{A}$., with elements given by

### v4

@@ I Introduction
The building blocks of communications and radar systems have been traditionally designed separately and assumed to utilize orthogonal resources. Integrated sensing and communications (ISAC)  [^1] Also commonly referred to as joint communication and sensing (JCAS) or joint communications and radar (JCR). [2, 3, 4, 5], a recently-emerged technology seen by many as a key driver of future mobile communication systems, challenges that established concept by considering the simultaneous provision of radar and communications functionalities under the same architecture, having the advantages of more efficient use of resources, lower power consumption, decreased hardware costs and the potential for wide market penetration [6, 7, 8]. While different perspectives on ISAC exists, each with a particular paradigm – $e.g.$ co-existence of communications and radar systems [9] or dual-function approaches [10] – the focus of this manuscript is on enabling sensing functionalities under wireless communications systems, also referred to as communication-centric ISAC.
Staying true to the integrative aspect of the proposition, earlier contributions to ISAC focused on schemes employing existing communications waveforms. In particular, many methods found in current literature seek to take advantage of certain features of channel estimation – especially in higher frequency bands where less scattering, higher resolution and domain-specific sparsity enable the estimation of individual channel paths – to incorporate radar-like functionalities, such as the estimation of radar parameters from echoes of the transmit signal [11, 12, 13, 14, 15].
Examples of the latter are techniques based on the orthogonal frequency division multiplexing (OFDM) waveform [13, 14] which, however, indicated that the high inter-carrier interference (ICI) can be a bottleneck that reduces robustness to high Doppler shifts present in doubly-dispersive environments, leading to severe degradation of performance. This factor alone, introduced by the channel diagonals spreading into a decaying band [16], severely hinders the use of OFDM in high mobility scenarios [17], as envisioned for 6G systems.
Another potential candidate for the ISAC paradigm is the orthogonal time frequency space (OTFS) waveform, which was originally proposed in [18] as a novel two-dimensional (2D) modulation scheme that embeds information directly on the delay-Doppler (DD) domain. Since the DD domain has a direct parallel to common radar processing framework, OTFS was quite popular at its inception, attracting much work that demonstrated its effectiveness for ISAC [11, 12, 13, 15, 14]. However, besides the significantly higher implementation complexity compared to OFDM, which is inherent to its 2D modulation approach, another limitation of OTFS is that it does not achieve the optimal diversity order in doubly-dispersive channels [19].
@@
It has been in fact demonstrated that AFDM has a superior bit error rate (BER) performance compared to OFDM, achieving also BERs comparable to OTFS while requiring lower overhead for channel estimation [22] as a result of the one-dimensional (1D) null-guard band required, as opposed to that of OTFS, which is 2D. In turn, evidence that AFDM can be effectively utilized for ISAC has also been given, for instance in [21] where a simple matched-filter (MF) approach was suggested. Building on the latter, an improved AFDM-based monostatic ISAC system with inherent adaptation to self-interference mitigation was also designed in [23], albeit at the expense of a significant increase in computational complexity due to the exhaustive maximum likelihood (ML) grid search required for the radar parameter estimation (RPE) procedure.
@@
One aspect of the ISAC concept that was often overlooked, and which recently has been addressed more frontally, however, is the impact of pilot overhead onto the overall approach, when designing the appropriate frame structure for optimum data decoding at a user equipment (UE) being served. It has been shown in [22], for instance, that a major advantage of AFDM over OFDM and OTFS is the lighter requirement in the number of pilots, which in turn can be taken advantage of by enabling the base station (BS) to perform RPE without expensive self-interference cancellation (SIC) techniques. The trend to reduce pilot overhead suggests, therefore, the adoption of joint channel and data estimation (JCDE) technology, which has been shown to be very effective in allowing efficient channel estimation with very few pilots [^2] The embedded channel estimation method proposed in [22] requires a single pilot, but surrounded by a null-guard interval, such that the fraction of the AFDM frame dedicated to payload is affected not just by the sole pilot but also the additional power that can be allocated to it. and under strenuous conditions [24, 25, 26, 27, 28, 1, 29], significantly reduce the overhead of coherent wireless systems. One example of the latter is the Bayesian bilinear inference (BBI) JCDE scheme for OFDM systems proposed in [28], which was shown to have the ability of handling channel aging and to entirely remove the need of pilots in favor of adequate coding at an equivalent rate. Another is the work in [1], where a Bayesian parametric bilinear Gaussian belief propagation (PBiGaBP) JCDE scheme for OTFS systems was proposed, which was shown to asymptotically approach the idealized scheme where perfect channel state information (CSI) knowledge is available.
@@
In view of all the above, we propose in this paper a combined and generalized JCDE-enabled ISAC solution for systems utilizing signals known to cope well with doubly-dispersive channels, including OFDM, OTFS and AFDM. The proposed solution comprises two main contributions. The first is a new JCDE mechanism, designed with basis on the Bayesian PBiGaBP originally proposed for OTFS waveforms in [1], which is shown to function also for the other modulation schemes under the general doubly-dispersive channel model developed in [16]. Comparisons of the state-of-the-art (SotA) OTFS and the contributed AFDM methods against a genie-aided linear Gaussian belief propagation (GaBP) benchmark scheme, in which perfect CSI is assumed, demonstrate that the proposed solution under AFDM achieves the same performance of the OTFS SotA, both of which also closely approach the idealistic benchmark, indicating that there is no penalty for the reduced modulation complexity advantage obtained when utilizing AFDM instead of OTFS, and that the proposed scheme is effective in reducing pilot overhead.
@@
The second contribution is a new RPE scheme based on the vectorized probabilistic data association (PDA) approach [25, 28] incorporating a Bernoulli-Gaussian denoiser tuned via expectation maximization (EM) [30] which quickly converges to the appropriate distribution parameters. Compared to the SotA alternative adapted from [31], which is designed around the sparse Bayesian learning (SBL) algorithm, the contributed method is shown to outperform the latter over any of the relevant waveforms, namely, OFDM, OTFS in [31] and AFDM. The technique also has an inherent and substantial computational cost advantage over ML-based SotA alternatives such as those in [22] and [23].
@@
  • By concisely summarizing the unified OFDM, OTFS and AFDM doubly-dispersive channel model from [16], such that message-passing receivers for such waveforms can be designed under a common structure, we propose in Section III a PBiGaBP-based scheme for JCDE, which applies to all the aforementioned waveforms.
  • We show that even if a single pilot per block is used, the proposed JCDE-AFDM system significantly outperforms both the OFDM and OTFS alternatives [1], as well as the AFDM method in [23], even when the latter is granted a substantial power advantage.
  • Finally, in Section IV-B, a novel PDA-based technique for RPE in doubly-dispersive channels is proposed, which in addition to enabling monostatic implementation for AFDM systems without requiring full-duplex (FD) technology, is shown to outperform the SotA of [31] for both OTFS and AFDM waveforms.
@@ II System Model
We follow [23] and consider an ISAC scenario composed of a single-antenna [^3] Both proposed algorithms can be extended straightforwardly to a single-input multiple-output (SIMO) case, with some modifications due to spatial filtering/interference cancellation for which isotropic beamforming as done in [32] could be used. Generalization to the multiple-input multiple-output (MIMO) case is also possible, but requires the formulation and solution of a non-trivial transmit (TX) and receive (RX) beamforming optimization problem. Such contribution falls outside the scope of this article and will be offered in a follow-up work. BS transmitting in downlink and taking advantage of first-order reflected (echo) signals to perform RPE in order to detect the ranges and velocities of objects (targets) in its surroundings, while the UE performs JCDE as illustrated in Figure 4. As depicted in the figure, it is assumed that a total of $P_{\mathrm{B}}$ target-originated echo signals reach the BS, which might be either from UE or other objects in the ambient. In turn, each UE is assumed to receive the signal transmitted by the BS scattered onto $P_{\mathrm{U}}+1$ paths, with $p_{\mathrm{U}}=0$ denoting the line-of-sight (LoS), without loss of generality (wlg), and the remaining $P_{\mathrm{U}}\leq P_{\mathrm{B}}$ paths being non-LoS (NLoS) components of the multi-path channel.
For the sake of simplicity and again wlg, we shall hereafter attain to a single user, since the multi-user (MU) case has no impact onto the JCDE procedure performed independently at each user, and is transparent to the RPE algorithm, as long as the radar receiver has sufficient degrees-of-freedom (DoF) to resolve all $P_{\mathrm{B}}$ echoed signals it receives. Also for simplicity, although we differentiate $P_{\mathrm{B}}$ and $P_{\mathrm{U}}$ in Figure 4, in the equations that follow, the number of paths/targets will be denoted by $P$, with the distinction between the two cases left to context.
$\mathbf{\Phi}_{p}\triangleq\text{diag}\bigg{(}[\overbrace{e^{-j2\pi\phi_{\mathrm{CP}}(\ell_{p})},e^{-j2\pi\phi_{\mathrm{CP}}(\ell_{p}-1)},\dots,e^{-j2\pi\phi_{\mathrm{CP}}(2)},e^{-j2\pi\phi_{\mathrm{CP}}(1)}}^{\ell_{p}\;\text{terms}},\overbrace{1,1,\dots,1,1}^{N-\ell_{p}\;\text{ones}}]\bigg{)}\in\mathbb{C}^{N\times N}.\vspace{-2ex}$ (5)
@@
Figure 1: Illustration of an ISAC system with $P_{\mathrm{B}}$ target echoes received by a single antenna [^4] Although it is possible to achieve $360^{\circ}$ coverage even in millimeter-wave (mmWave) bands with antenna arrays or omnidirectional antennas [33, 34, 35], for which Figure 4 would apply as is, this illustration is not meant to be restricted to any specific band nor should it be interpreted as exclusive to the omnidirectional scenario, but rather can also be envisioned for sectorized scenarios as well [11].BS, and $P_{\mathrm{U}}$ resolvable paths seen at the UE.
@@
II-A Generalized Doubly-Dispersive Channel Model
@@
where $t$ and $\tau$ are the continuous time and delays, respectively.
Leveraging equation (1), the input-output relationship of the channel in the TD can be obtained via the linear convolution of the transmit signal $s(t)$ and channel impulse response as described in [38, eq. (10.69)], which added with noise yields
$r(t)=s(t)*h(t,\tau)+w(t)$
@@
Defining $r[n]$ and $s[n]$, with $n\in\{0,\dots,N-1\}$, to be the sampled sequences of $r(t)$ and $s(t)$, respectively, with samples taken at a sufficiently high sampling rate $f_{\mathrm{S}}\triangleq\frac{1}{T_{\mathrm{S}}}$[Hz], the following discrete equivalent of equation (II-A) can be obtained
$r[n]=\sum_{\ell=0}^{\infty}s[n-\ell]\bigg{(}\sum_{p=0}^{P}h_{p}\cdot e^{j2\pi\frac{\nu_{p}}{f_{\mathrm{S}}}n}\cdot\delta\Big{(}\ell-\frac{\tau_{p}}{T_{\mathrm{S}}}\Big{)}\bigg{)}+w[n],\vspace{-1ex}$ (3)
where $\ell$ is the normalized discrete delay index, and $T_{\mathrm{S}}$ the delay resolution as well as the sampling interval, such that $\tau=\ell\cdot T_{\mathrm{S}}$.
Finally, defining the normalized digital Doppler shift and the normalized discrete delay of the $p$-th path respectively as $f_{p}\triangleq\frac{N\nu_{p}}{f_{\mathrm{S}}}$ and $\ell_{p}\triangleq\frac{\tau_{p}}{T_{\mathrm{S}}}$, where the sampling rate $f_{\mathrm{S}}$ is chosen sufficiently high (via oversampling [16] if necessary) [^6] The extension to incorporate normalized fractional delays if the chosen $f_{\mathrm{S}}$ is insufficient to render the integer assumption of the normalized delays is trivial and can be done via the approach suggested in [39]. Subsequently, since the only difference would be a modification to $\mathbf{\Pi}$, both proposed algorithms could still be utilized as they are with no changes. to ensure that $\ell_{p}-\lfloor\frac{\tau_{p}}{T_{\mathrm{S}}}\rceil\approx 0$, the discrete convolution described in equation (3) can be rewritten in terms of the circular convolution
$\mathbf{r}=\bigg{(}\sum_{p=0}^{P}h_{p}\cdot\mathbf{\Phi}_{p}\cdot\bm{\Omega}^{f_{p}}\cdot\mathbf{\Pi}^{\ell_{p}}\bigg{)}\cdot\mathbf{s}+\mathbf{w}=\mathbf{H}\cdot\mathbf{s}+\mathbf{w},\vspace{-1ex}$ (4)
where the inclusion and posterior removal of a cyclic prefix (CP) of length $N_{\mathrm{CP}}$ to the sequence $s[n]$ is already taken into account [16].
In equation (4), $\mathbf{r}\triangleq[r[0],\cdots,r[N-1]]\in\mathbb{C}^{N\times 1}$ and $\mathbf{s}\triangleq[s[0],\cdots,s[N-1]]\in\mathbb{C}^{N\times 1}$ are the transmit and received signal sample vectors; $\mathbf{w}\in\mathbb{C}^{N\times 1}$ denotes AWGN; $\mathbf{H}\in\mathbb{C}^{N\times N}$ is the effective channel matrix; $\mathbf{\Phi}_{p}\in\mathbb{C}^{N\times N}$ described in equation (5) is a diagonal matrix which captures the effect of the CP onto the $p$-th channel path, where $\phi_{\mathrm{CP}}(n)$ denotes a phase function on the sample index $n\in\{0,\cdots,N-1\}$ [16] depending on the specific waveform used; $\bm{\Omega}\in\mathbb{C}^{N\times N}$, described in equation (6), is a diagonal matrix containing the $N$ complex roots of one and $\mathbf{\Pi}\in\{0,1\}^{N\times N}$ is the forward cyclic shift matrix [^7] To clarify, the matrix $\mathbf{\Pi}$ is such that $\bm{A}\cdot\mathbf{\Pi}^{\ell_{p}}$, with $\ell_{p}\in\mathbb{N}_{0}$, is a cyclic left-shifted version of the matrix $\bm{A}$, such that the first $\ell_{p}$ columns of $\bm{A}$ are moved to the positions of the last $\ell_{p}$ columns. Notice also that $\mathbf{\Pi}^{0}$ is the $N\times N$ identity matrix, such that $\bm{A}\cdot\mathbf{\Pi}^{0}=\bm{A}$., with elements given by
@@
To that end, we consider the OFDM, OTFS and AFDM waveforms, which are known to perform well in doubly-dispersive channels. Since a thorough comparison of the different individual features of each of them has been done in [16], we limit ourselves here to briefly revise the models introduced thereby with the objective of obtaining a general mathematical formulation for the design of receivers/estimators for these waveforms under a common framework.